How to find limiting reagent? i am given 4.3mL of isopentanol and acetic acid (4eq)
1
Expert's answer
2018-10-01T08:02:02-0400
Isopentanol reacts with acetic acid to form ester.
C5H11OH + CH3COOH = CH3COOC5H11 + H2O
V(C5H11OH) = 4.3 mL
density of isopentanol = 0.81 g/mL
m(C5H11OH) = 0.81g/mL*4.3mL = 3.483 g
M(C5H11OH) = 88 g/mol
n(C5H11OH) = m/M = 3.483g/88g/mol = 0.04 mol
n(CH3COOH) = 4 mol , because acetic acid is monobasic acid.
According to the equation 1 mol of isopentanol reacts with 1 mol of acetic acid, so isopentanol is the limiting reagent, acetic acid is in the excess.